Johnny is designing a rectangular poster to contain 48inches squared of printing with a 4-in margin at the top and bottom and a 3-in margin at each side. What overall dimensions will minimize the amount of paper​ used?


1
Expert's answer
2021-10-13T17:28:44-0400

Let the dimensions of the poster be x by y inches.

"xy=48\to y=\frac{48}{x}."

"A=(x+6*2)(y+8*2)=xy+16x+12y+192=16x+\frac{576}{x}+240."

"\frac{dA}{dx}=0\to16-\frac{576}{x^2}=0\to x^2=36."

"x=\sqrt{36}=6."

"y=\frac{48}{6}=8."

"x+6*2=6+12=18."

"y+8*2=8+16=24."

The overall dimensions of the paper: 18 by 24 inches.